A passenger train travels 150 miles in the same time that a frieght train travels 100 miles. If the frieght train travels 20 mph slower than the passenger train, what is the speed of the passenger train?
Answer by stanbon(48558)

A passenger train travels 150 miles in the same time that a frieght train travels 100 miles. If the frieght train travels 20 mph slower than the passenger train, what is the speed of the passenger train?
----------------
Passenger train DATA:
distance = 150 mi ; Rate = x mph ; time = d/r = 150/x hrs
------------------
Freight train DATA:
disance = 100 mi ; Rate = x-20 mph ; time = d/r = 100/(x-20) hrs
---------------
EQUATION:
time = time
150/x = 100/(x-20)
3/x = 2/(x-20)
3x-60 = 2x
x=60 mph (passenger train rate)
x-20 = 40 mph (freight train rate)
================
Cheers,
Stan H.